OpenStudy (anonymous):

Lisa's piggy bank contained 21 dimes, nickels, and pennies. The total in the piggy bank was $1.21. Lisa removed all the dimes from the piggy bank. This left only 41 cents in the bank. How many pennies were in the bank?

OpenStudy (anonymous):

x = number of dimes y = number of nickels z = number of pennies x + y + z = 21 (.1)*x + (.05)*y + (.01)*z = 1.21 (.05)*y + (.01)*z = .41 Take equation 3 and solve for y: (.05)*y = .41 - (.01)*z y = 8.2 - (.2)*z Plug this value for y into equation 1 and solve for x: x + (8.2 - (.2)*z) + z = 21 x + z - (.2)*z = 21 - 8.2 x + (.8)*z = 12.8 x = 12.8 - (.8)*z Now, plug your values for x and y into equation 2 and solve for z: (.1)*(12.8 - (.8)*z) + (.05)*(8.2 - (.2)*z) + (.01)*z = 1.21 1.28 - (.08)*z + .41 - (.01)*z + (.01)*z = 1.21 1.69 - (.08)*z = 1.21 (.08)*z = .48 z = 6 Now, plug your value for z back into the simplified equation for y: y = 8.2 - (.2)*(6) y = 7 So, Lisa had 7 nickels and 6 pennies in the bank.

OpenStudy (anonymous):

That is a lot of work!..

OpenStudy (anonymous):

im really bored lol

OpenStudy (anonymous):

Oh.